In 2000 the population density of the United States was 80 people per square mile, and in 2011 it was 88 people per square mile. Use a linear function to estimate when the U.S. population will reach 91 people per square mile.

Answer by ewatrrr(24785)

2011 , 88
2000 , 80
m = 8/11
y -80 = (8/11)(x - 2000)
y = (8/11)x -(8/11)2000 + 80
|
91 = (8/11)x -(8/11)2000 + 80
91 +(8/11)2000 - 80= (8/11)x
(11/8)[11 + (8/11)2000 ] = x
121/8 + 2000 = x = 2015.125
In the year 2015